Mean and Variance of Sampling Method:


  1. What is the summation of x squared if the given are the following: 50, 46, 51, 45, 40, 43, 57, 58, 64, 59, 60, 62? ∑𝑥2?
  2. What is the summation of x if the given are the following: 50, 46, 51, 45, 40, 43, 57, 58, 64, 59, 60, 62? ∑𝑥?
  3. What is the summation of the quantity of x minus the mean close quantity squared if the given are the following: 100; 95; 98; 97; 93; 92; 91; 96; 99; 89? ∑(𝑥−𝑥̅ )2?
  4. What is the variance if the given are the following: 100; 95; 98; 97; 93; 92; 91; 96; 99; 89?
  5. What is the standard deviation if the given are the following: 100; 95; 98; 97; 93; 92; 91; 96; 99; 89?
  6. What is the summation of the quantity of x minus the mean close quantity squared if the given are the following: 50, 46, 51, 45, 40, 43, 57, 58, 64, 59, 60, 62? ∑(𝑥− 𝑥̅ )2?
1
Expert's answer
2022-03-30T08:31:42-0400

1."\sum x^2=2500+2116+2601+2025+1600+1849+3249+3364+4096+3481+3600+3844=34325"

2.

"\sum x=50+46+51+45+40+43+57+58+64+59+60+62=636"

3. "\mu=(100+95+98+97+93+92+91+96+99+89)/10=950/10=95"

"\sum (x-\mu)^2=25+0+9+4+4+9+16+1+16+36=120"

4. "\sigma^2=\frac{ \sum (x-\mu)^2}{n-1}=120/9=13.3"

5. "\sigma=\sqrt{\sigma^2}=3.65"

6. "\u200b\n\\sum (x-\\mu)^2=9+49+4+64+169+100+16+25+121+36+49+81\n\u200b\n =723"